OpenStudy (anonymous):

what is half of 2 7/8 ?

OpenStudy (anonymous):

What is half of 2 7/8 1. halve the integer to give 1; halve the fraction to give 7/16. Put them together to give 1 7/16. 2.2 7/8 = 23/8 x 1/2 = 23/16 = 1 7/16

OpenStudy (anonymous):

1-7/16

OpenStudy (anonymous):

1.79